OpenStudy (anonymous):

the scores on an exam are normally distributed, with a mean of 74 and a standard deviation of 7. what percent of the scores are less than 81?

jimthompson5910 (jim_thompson5910):

Hint: notice how 81 is 1 standard deviation above the mean (74+7 = 81)

OpenStudy (anonymous):

I still have no clue..... SO LOST

jimthompson5910 (jim_thompson5910):

do you remember the empirical rule?

OpenStudy (kagıtucak):

\[PDF of this variable=1/(2\pi7^2)\exp[-(x-74)^2/(2.7^2)]\] But the question asks the CDF of this varible when it is equal to 81 and \[FX(x)=\int\limits_{-infinity}^{x}fX(x)dx\]

OpenStudy (kagıtucak):

you need to take integral of the pdf and put 81 to into it